# Statistical Finite Elements via Interacting Particle Langevin Dynamics

### v1

@@ 3 Estimating Forcing via Interacting Particle Langevin Dynamics
$\mathbf{u}_{k+1}^{(n)}$ $=\mathbf{u}_{k}^{(n)}-\gamma\nabla_{\mathbf{u}}\Phi^{y}(\mathbf{u}_{k}^{(n)},\mathbf{b}_{k})+\sqrt{2\gamma}\,\bm{\zeta}_{k+1}^{(n)},\quad\forall n\in\left\{1,\dots,N\right\}.$ (22)
We consider linear and nonlinear examples, deriving the form of these potentials in both cases.
@@ 3.1 Forcing Estimation for Linear Poisson PDE
One example of an elliptic PDE is the Poisson equation, which describes processes such as steady-state heat diffusion. The strong form of the linear Poisson equation with homogeneous diffusivity coefficient equal to one, is given by
$-\nabla\cdot(\nabla u(x))$ $=f+\epsilon,\quad x\in\Omega$ (23)
$u(x)$ $=0,\quad x\in\partial\Omega.$ (24)
The bilinear form associated with this strong form, $\mathcal{A}(u,v)=\left\langle\nabla u,\nabla v\right\rangle$, produces the so called stiffness matrix, $\left[\mathbf{A}\right]_{ij}=\left\langle\nabla\phi_{i},\nabla\phi_{j}\right\rangle$. The stiffness matrix is symmetric positive definite meaning the linear system, $\mathbf{A}\mathbf{u}=\mathbf{b}$, will always have a unique solution and the inverse of the stiffness matrix in (10) is well defined. The full joint data generating model factorises as the product of the following Gaussian distributions
$p(\mathbf{b})=\mathcal{N}\left(\bm{\mu},\bm{\Sigma}\right),\quad p(\mathbf{u}|\theta,\mathbf{b})=\mathcal{N}\left(\mathbf{A}^{-1}\mathbf{b},\mathbf{A}^{-1}\mathbf{G}\mathbf{A}^{-\top}\right),\quad p(\mathbf{y}|\mathbf{u})=\mathcal{N}\left(\mathbf{H}\mathbf{u},\mathbf{R}\right),$ (25)
@@
where the discretised prior distribution can be constructed directly from a GP prior over the forcing function, i.e. given $f\sim\mathcal{GP}(\mu_{f},k_{f})$, we have $\left[\bm{\mu}\right]_{i}=\langle\phi_{i},\mu_{f}\rangle$, $\left[\bm{\Sigma}\right]_{ij}=\langle\phi_{i},\langle k_{f}(\cdot,\cdot),\phi_{j}\rangle\rangle$. We also apply a fixed preconditioner to this potential to improve stability and sample efficiency. The preconditioners for the latent sampling $\mathbf{P}_{\mathbf{u}}$, and for parameter gradient descent $\mathbf{P}_{\mathbf{b}}$ are set to the inverse of the block diagonal Hessian matrices (see Equation (40)) as follows:
$\mathbf{P}_{\mathbf{b}}=\left[\mathbf{G}^{-1}+\bm{\Sigma}^{-1}\right]^{-1},\quad\mathbf{P}_{\mathbf{u}}=\left[\mathbf{A}^{\top}\mathbf{G}^{-1}\mathbf{A}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right]^{-1},$ (26)
@@
which individually preconditions both the posterior sampling and the parameter gradient descent with their exact Hessians, but ignores their interaction (off-diagonal terms of the full block Hessian). With this choice of preconditioning, the forcing estimation simplifies into the following update steps
$\mathbf{b}_{k+1}$ $=(1-\gamma)\mathbf{b}_{k}+\gamma\mathbf{P}_{\mathbf{b}}\left(\mathbf{G}^{-1}\mathbf{A}\left(\frac{1}{N}\sum_{n=1}^{N}\mathbf{u}_{k}^{(n)}\right)+\bm{\Sigma}^{-1}\bm{\mu}\right)+\sqrt{\frac{2\gamma}{N}}\mathbf{P}_{\mathbf{b}}^{1/2}\bm{\zeta}_{k+1}^{(0)},$ (27)
$\mathbf{u}_{k+1}^{(n)}$ $=(1-\gamma)\mathbf{u}_{k}^{(n)}+\gamma\mathbf{P}_{\mathbf{u}}\left(\mathbf{A}^{\top}\mathbf{G}^{-1}\mathbf{b}_{k}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{y}\right)+\sqrt{2\gamma}\mathbf{P}_{\mathbf{u}}^{1/2}\bm{\zeta}_{k+1}^{(n)},$ (28)
for all $n\in\left\{1,\dots,N\right\}$.
3.2 Forcing Estimation for Nonlinear Poisson PDE
$-\nabla\cdot(q(u)\nabla u)=f+\epsilon.$ (29)
For full details see Appendix C. With a basis function expansion $u_{h}=\sum_{i=1}^{n_{u}}\hat{u}_{i}\phi_{i}$, this is rewritten in vector form
where we have $\mathbf{u}=\left[\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{n_{u}}\right]^{\top}$, $\left[\mathbf{F}(\mathbf{u})\right]_{i}=\left\langle q(u_{h})\nabla u_{h},\nabla\phi_{i}\right\rangle$, $\left[\mathbf{b}\right]_{i}=\langle f,\phi_{i}\rangle$, and $\left[\mathbf{G}\right]_{ij}=\langle\phi_{i},\langle k(\cdot,\cdot),\phi_{j}\rangle\rangle$. In this case the computation of the gradient of the potential with respect to $\mathbf{u}$, $\nabla_{\mathbf{u}}\Phi^{y}(\mathbf{u},\mathbf{b})$ is intractable due to the log-determinant Jacobian term in the score of the nonlinear statFEM prior, so we turn to approximations of this distribution to make progress. Approximating this distribution with a Gaussian alleviates the problem of intractability, providing a linear score computed using the approximate mean and covariance. We present here the algorithm for sampling based on a Gaussian approximation of the nonlinear statFEM prior $p(\mathbf{u}|\theta,\mathbf{b})\approx\mathcal{N}(\mathbf{m},\mathbf{C})$:
$\mathbf{b}_{k+1}$ $=(1-\gamma)\mathbf{b}_{k}+\gamma\mathbf{P}_{\mathbf{b}}\left(\mathbf{G}^{-1}\left(\frac{1}{N}\sum_{n=1}^{N}\mathbf{F}(\mathbf{u}_{k}^{(n)})\right)+\bm{\Sigma}^{-1}\bm{\mu}\right)+\sqrt{\frac{2\gamma}{N}}\mathbf{P}_{\mathbf{b}}^{1/2}\zeta_{k+1}^{(0)}$
$\mathbf{u}_{k+1}^{(n)}$ $=(1-\gamma)\mathbf{u}_{k}^{(n)}+\gamma\mathbf{P}_{\mathbf{u}}\left(\mathbf{C}^{-1}\mathbf{m}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{y}\right)+\sqrt{2\gamma}\mathbf{P}_{\mathbf{u}}^{1/2}\zeta_{k+1}^{(n)},\quad\forall n\in\left\{1,\dots,N\right\}$
$\text{with}\quad\mathbf{P}_{\mathbf{b}}=\left[\mathbf{G}^{-1}+\bm{\Sigma}^{-1}\right]^{-1},\quad\mathbf{P}_{\mathbf{u}}=\left[\mathbf{C}^{-1}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right]^{-1}.$
Note the gradient with respect to the forcing can be computed exactly because the potential is quadratic in $\mathbf{b}$, i.e. $\nabla_{\mathbf{b}}\Phi^{y}(\mathbf{u},\mathbf{b})=\mathbf{G}^{-1}\left(\mathbf{F}(\mathbf{u})-\mathbf{b}\right)$. Simple approaches to approximating the distribution of nonlinearly transformed Gaussian random variables [25] include a first-order Taylor approximation, the unscented transform, and an empirical approximation to the first two moments with a sample based approach (See Table 1).
Table 1: Gaussian approximations to nonlinearly transformed Gaussian random variables: first-order Taylor (FOT), unscented transform (UT) and Monte Carlo (MC). FOT requires inversion of the Jacobian, $\mathbf{J}(\hat{\mathbf{m}})=\frac{\partial\mathbf{F}}{\partial\mathbf{u}}|_{\hat{\mathbf{m}}}$. UT requires $12n_{u}+$ 1 solves, details in Appendix D.2. MC requires $M$ solves, where $\tilde{\mathbf{b}}^{(j)}\sim\mathcal{N}(\mathbf{b},\mathbf{G})$.
Solve Mean: $\mathbf{m}$ Covariance: $\mathbf{C}$
Details of each of these approximations are given in Appendix D.2.
4 Solving the Inverse Problem via Interacting Particle Langevin Dynamics
In this section, we focus on the inverse problem of estimating $\theta$ for fixed $\mathbf{b}$, and we place a prior on $\theta$, hence we again introduce a simplified notation:
$\Psi_{\theta}(\mathbf{u}):=-\log p(\mathbf{u}|\theta,\mathbf{b})-\log p(\theta),\quad\Psi_{\theta}^{y}(\mathbf{u}):=\Psi_{\theta}(\mathbf{u})-\log p(\mathbf{y}|\mathbf{u}),$ (32)
where $\Psi_{\theta}(\mathbf{u}),\Psi_{\theta}^{y}(\mathbf{u})$ are the prior and posterior potentials respectively. The general form of the IPLA iterations for solving the inverse problem are given below
$\theta_{k+1}$ $=\theta_{k}-\frac{\gamma}{N}\sum_{n=1}^{N}\nabla_{\theta}\Psi_{\theta_{k}}^{y}(\mathbf{u}_{k}^{(n)})+\sqrt{\frac{2\gamma}{N}}\,\bm{\zeta}_{k+1}^{(0)}$ (33)
$\mathbf{u}_{k+1}^{(n)}$ $=\mathbf{u}_{k}^{(n)}-\gamma\nabla_{\mathbf{u}}\Psi_{\theta_{k}}^{y}(\mathbf{u}_{k}^{(n)})+\sqrt{2\gamma}\,\bm{\zeta}_{k+1}^{(n)},$ (34)
for $n\in\left\{1,\dots,N\right\}$. We now will apply this algorithm to the problem of estimating the diffusivity parameter in the linear Poisson equation.
4.1 Estimating Diffusivity Parameter for Linear Poisson PDE
The linear Poisson example in section 3.1 assumes a known, fixed diffusivity. Here, we wish to estimate this diffusivity using our algorithm, assuming it is unknown a priori. Introducing this as a parameter, $\varkappa$, we have the strong form of the linear Poisson equation with homogeneous diffusivity
$-\nabla\cdot(\varkappa\nabla u)=f+\epsilon.$ (35)
Estimating this homogeneous diffusivity term, in practice, requires a mapping that enforces non-negativity to ensure the parameter gradient descent produces a physically valid system. In our case we use the exponential mapping to reparametrise the strong form with diffusivity $\varkappa$, by $\theta=\log\varkappa$ to ensure non-negativity. The data generative model follows
$p(\theta)=\mathcal{N}\left(\mu_{\theta},\sigma^{2}_{\theta}\right),\quad p(\mathbf{u}|\theta,\mathbf{b})=\mathcal{N}\left(e^{-\theta}\mathbf{A}^{-1}\mathbf{b},e^{-2\theta}\mathbf{A}^{-1}\mathbf{G}\mathbf{A}^{-\top}\right),\quad p(\mathbf{y}|\mathbf{u})=\mathcal{N}\left(\mathbf{H}\mathbf{u},\mathbf{R}\right).$
This is equivalent to a log-normal prior over the diffusivity, but with the added constraint of non-negativity when performing gradient update steps. Preconditioning is applied at each time-step to the particle Langevin dynamics, based on the current value of $\theta$, where $\mathbf{P}_{\mathbf{u}}[\theta]=\left[e^{2\theta}\mathbf{A}^{\top}\mathbf{G}^{-1}\mathbf{A}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right]^{-1}$. Writing out the IPLA algorithm for this problem, we have
@@ 4.1 Estimating Diffusivity Parameter for Linear Poisson PDE
$\theta_{k+1}$ $=\theta_{k}-\gamma\left(n_{u}+\frac{\theta_{k}-\mu_{\theta}}{\sigma_{\theta}^{2}}\right)-\frac{\gamma}{N}\sum_{n=1}^{N}e^{\theta_{k}}\left(e^{\theta_{k}}\mathbf{A}\mathbf{u}_{k}^{(n)}-\mathbf{b}\right)^{\top}\mathbf{G}^{-1}\mathbf{A}\mathbf{u}_{k}^{(n)}+\sqrt{\frac{2\gamma}{N}}\,\bm{\zeta}_{k+1}^{(0)}$ (36)
@@
$\mathbf{u}_{k+1}^{(n)}$ $=(1-\gamma)\mathbf{u}_{k}^{(n)}+\gamma\mathbf{P}_{\mathbf{u}}[\theta_{k}]\left(e^{\theta_{k}}\mathbf{A}^{\top}\mathbf{G}^{-1}\mathbf{b}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{y}\right)+\sqrt{2\gamma}\,\mathbf{P}_{\mathbf{u}}[\theta_{k}]^{1/2}\bm{\zeta}_{k+1}^{(n)}$ (37)
for all $n\in\left\{1,\dots,N\right\}$.
@@ 5.1 Convergence of elliptic PDE forcing estimate to MMAP estimator
Recall that we consider the setting in section 3.1 and in this case, we have the potential $\Phi(\mathbf{u},\mathbf{b})$ for the joint system $\mathbf{v}=(\mathbf{b},\mathbf{u})$ conditioned on the data $\mathbf{y}$. We can write the gradients of the potential w.r.t. the solution, $\mathbf{u}$, and parameters $\mathbf{b}$, are as follows
$\nabla_{\mathbf{u}}\Phi^{y}(\mathbf{u},\mathbf{b})$ $=\mathbf{A}^{\top}\mathbf{G}^{-1}\left(\mathbf{A}\mathbf{u}-\mathbf{b}\right)+\mathbf{H}^{\top}\mathbf{R}^{-1}\left(\mathbf{H}\mathbf{u}-\mathbf{y}\right),$ (38)
$\nabla_{\mathbf{b}}\Phi^{y}(\mathbf{u},\mathbf{b})$ $=-\mathbf{G}^{-1}\left(\mathbf{A}\mathbf{u}-\mathbf{b}\right)+\bm{\Sigma}^{-1}(\mathbf{b}-\bm{\mu}).$ (39)
For this problem we have a quadratic potential, and so the strong convexity and Lipschitz constants are determined by the minimum and maximum eigenvalues of the Hessian. We show first below that the potential is strongly convex and its gradient is Lipschitz continuous.
@@
The potential $\Phi^{y}(\mathbf{u},\mathbf{b})$ is $\mu$-strongly convex, with $\mu=\lambda_{\mathrm{min}}\left(\nabla^{2}\Phi^{y}\right)$ and the gradient of the potential $\Phi^{y}(\mathbf{u},\mathbf{b})$ is $L$-Lipschitz, with $L=\lambda_{\mathrm{max}}\left(\nabla^{2}\Phi^{y}\right)$.
Proof.
The strong convexity constant follows simply from
$\left\langle\mathbf{v}-\mathbf{v}^{\prime},\nabla\Phi^{y}(\mathbf{v})-\nabla\Phi^{y}(\mathbf{v}^{\prime})\right\rangle$ $=\left(\mathbf{v}-\mathbf{v}^{\prime}\right)^{\top}\underbrace{\begin{bmatrix}\mathbf{A}^{\top}\mathbf{G}^{-1}\mathbf{A}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}&-\mathbf{A}^{\top}\mathbf{G}^{-1}\\ -\mathbf{G}^{-1}\mathbf{A}&\mathbf{G}^{-1}+\bm{\Sigma}^{-1}\end{bmatrix}}_{\nabla^{2}\Phi^{y}}\left(\mathbf{v}-\mathbf{v}^{\prime}\right)$
$\geq\mu\|\mathbf{v}-\mathbf{v}^{\prime}\|^{2},\quad\mu=\lambda_{\mathrm{min}}\left(\nabla^{2}\Phi^{y}\right)$ (40)
where $\lambda_{\mathrm{min}}\left(\cdot\right)$ denotes the minimum eigenvalue of the matrix. For the Lipschitz constant,
$\left\|\nabla\Phi^{y}(\mathbf{v})-\nabla\Phi^{y}(\mathbf{v}^{\prime})\right\|$ $=\left\|\nabla^{2}\Phi^{y}\left(\mathbf{v}-\mathbf{v}^{\prime}\right)\right\|$
$\leq L\|\mathbf{v}-\mathbf{v}^{\prime}\|,\quad L=\lambda_{\mathrm{max}}\left(\nabla^{2}\Phi^{y}\right)$ (41)
where $\lambda_{\mathrm{max}}\left(\cdot\right)$ denotes the maximum eigenvalue of the matrix. ∎
Provided the Hessian is positive-definite, we have $\mu>0$, and a unique minimiser. For the forcing estimation of an elliptic PDE, this can be found analytically and is derived in Appendix B. We experimentally confirm this result, by estimating the order of convergence of the IPLA estimate to the MMAP estimator in section 6.1.
Theorem 1.
Following Lemma 1, the parameter marginal $(\mathbf{b}_{k})_{k\geq 1}$ generated by the IPLA algorithm (27) (with identity preconditioners) and $\gamma\leq 2/(\mu+L)$ for estimating the forcing function in the linear Poisson equation converges to the MMAP estimator $\mathbf{b}^{\star}=\arg\max p(\mathbf{b}|\mathbf{y})$. In particular, we have the following convergence result
where $\nu_{0}$ is the initial distribution of the particle system so that $W_{2}(\nu_{0},\tilde{\pi})<\infty$ where $\tilde{\pi}$ is given in the proof, $\mu$ and $L$ is the strong convexity constant and the Lipschitz constant of the potential $\Phi^{y}(\mathbf{u},\mathbf{b})$, respectively, given in Lemma 1.
@@
A few remarks to fully unpack this result are in order.
@@
Remark 1 makes it clear that the number of iterations $k$ to obtain $\varepsilon$-accuracy is strongly influenced by $\kappa$. It is thus crucial to control this number for ill-conditioned problems, as in our case. The algorithm in (27) is preconditioned and this can potentially improve our bound. Specifically, with a transformation of variables $\mathbf{w}\coloneqq\mathbf{P}^{-1/2}\mathbf{v}$, it can be shown via Itô’s formula that an equivalent SDE, with potential $\Phi^{y}_{\mathbf{P}}(\mathbf{w})\coloneqq\Phi^{y}(\mathbf{v})=\Phi^{y}(\mathbf{P}^{1/2}\mathbf{w})$ has the same invariant measure [4]. We can then analyse the Lipschitz and strong convexity constants in the transformed coordinate system, using the same method as in section 5.1 now that the SDE is of the same form. The Hessian for the transformed potential $\nabla_{\mathbf{w}}^{2}\Phi^{y}_{\mathbf{P}}(\mathbf{w})=\mathbf{P}^{1/2}\nabla_{\mathbf{v}}^{2}\Phi^{y}(\mathbf{v})\mathbf{P}^{1/2}$, and so the constants relating to the preconditioned system are
$\mu_{\mathbf{P}}=\lambda_{\mathrm{min}}\left(\mathbf{P}^{1/2}\nabla^{2}\Phi^{y}\mathbf{P}^{1/2}\right),\quad L_{\mathbf{P}}=\lambda_{\mathrm{max}}\left(\mathbf{P}^{1/2}\nabla^{2}\Phi^{y}\mathbf{P}^{1/2}\right).$ (43)
The bound in Theorem 1 then applies directly with the constants $\mu_{\mathbf{P}}$ and $L_{\mathbf{P}}$. Let $\kappa_{\mathbf{P}}=L_{\mathbf{P}}/\mu_{\mathbf{P}}$ be the condition number of the preconditioned system. Given Remark 1 and assuming that $\kappa_{\mathbf{P}}\ll\kappa$, preconditioning has two clear effects: First, the step-size $\gamma=\mathcal{O}(\varepsilon^{2}n_{u}^{-1}\kappa_{\mathbf{P}}^{-2})$ can now be bigger (within the restrictions provided in Theorem 1). Secondly, to obtain $\varepsilon$-accuracy, we may require significantly fewer steps, i.e., $k=\widetilde{\mathcal{O}}(n_{u}\kappa_{\mathbf{P}}\varepsilon^{-2})$. $\diamond$

### v2

@@ 3 Interacting Particle Langevin Dynamics for Linear Poisson PDE
$\mathrm{d}\mathbf{u}_{t}^{(n)}$ $=-\nabla_{\mathbf{u}}\Psi_{\theta}^{y}(\mathbf{u}_{t}^{(n)},\mathbf{b}_{t})\mathrm{d}t+\sqrt{2}\,\mathrm{d}\mathbf{B}_{t}^{(n)},\quad\forall n\in\{1,\dots,N\},$ (23)
where $(\mathbf{B}_{t}^{(n)})_{t\geq 0}$ is an independent Brownian motion for each $n$. This can be discretised to form the update steps, which are given below for the kth iteration
$\mathbf{b}_{k+1}$ $=\mathbf{b}_{k}-\frac{\gamma}{N}\sum_{n=1}^{N}\nabla_{\mathbf{b}}\Psi_{\theta}^{y}(\mathbf{u}_{k}^{(n)},\mathbf{b}_{k})+\sqrt{\frac{2\gamma}{N}}\,\bm{\zeta}_{k+1}^{(0)}$ (24)
$\mathbf{u}_{k+1}^{(n)}$ $=\mathbf{u}_{k}^{(n)}-\gamma\nabla_{\mathbf{u}}\Psi_{\theta}^{y}(\mathbf{u}_{k}^{(n)},\mathbf{b}_{k})+\sqrt{2\gamma}\,\bm{\zeta}_{k+1}^{(n)},\quad\forall n\in\left\{1,\dots,N\right\},$ (25)
where for each $n\in\{1,\dots,N\}$, $\{\bm{\zeta}_{k}^{(n)}\}_{k\in\mathbb{N}}$ is a sequence of i.i.d. standard Gaussian random variables. This can simply be adapted to MMAP estimation of the diffusivity, by instead fixing $\mathbf{b}$, and replacing (22) to evolve $\theta$ with the gradient of the joint potential $\nabla_{\theta}\Psi_{\theta}^{y}(\mathbf{u},\mathbf{b})$. In the following two sections we derive the particular forms of the IPLA update steps for both the forcing estimation and solving the inverse problem, including preconditioning.
@@ 3.1 Forcing Estimation for Linear Poisson PDE
Here, we derive the IPLA update steps for estimating the forcing $\mathbf{b}$, while taking the diffusivity as known and constant. We also apply a fixed preconditioner to this potential to improve stability and sample efficiency. The preconditioners for the latent sampling $\mathbf{P}_{\mathbf{u}}$, and for parameter gradient descent $\mathbf{P}_{\mathbf{b}}$ are set to the inverse of the block diagonal Hessian matrices (see Equation (33)) as follows:
$\mathbf{P}_{\mathbf{b}}=\left[\mathbf{G}^{-1}+\bm{\Sigma}^{-1}\right]^{-1},\quad\mathbf{P}_{\mathbf{u}}=\left[\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right]^{-1},$ (26)
@@
which individually preconditions both the posterior sampling and the parameter gradient descent with their exact Hessians, but ignores their interaction (off-diagonal terms of the full block Hessian). With this choice of preconditioning the forcing estimation simplifies into the following update steps
@@
$\mathbf{b}_{k+1}$ $=(1-\gamma)\mathbf{b}_{k}+\gamma\mathbf{P}_{\mathbf{b}}\left(\mathbf{G}^{-1}\mathbf{A}_{\theta}\left(\frac{1}{N}\sum_{n=1}^{N}\mathbf{u}_{k}^{(n)}\right)+\bm{\Sigma}^{-1}\bm{\mu}\right)+\sqrt{\frac{2\gamma}{N}}\mathbf{P}_{\mathbf{b}}^{1/2}\bm{\zeta}_{k+1}^{(0)},$ (27)
$\mathbf{u}_{k+1}^{(n)}$ $=(1-\gamma)\mathbf{u}_{k}^{(n)}+\gamma\mathbf{P}_{\mathbf{u}}\left(\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\mathbf{b}_{k}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{y}\right)+\sqrt{2\gamma}\mathbf{P}_{\mathbf{u}}^{1/2}\bm{\zeta}_{k+1}^{(n)},$ (28)
for all $n\in\{1,\dots,N\}$, with $\{\bm{\zeta}_{k}^{(n)}\}_{n=0}^{N}$ as i.i.d. standard $n_{u}$-dimensional Gaussians.
3.2 Solving the Inverse Problem for Linear Poisson PDE
Here, we derive the update steps for estimating the diffusivity, given a known forcing function. Preconditioning is applied at each time-step to the particle Langevin dynamics, based on the current value of $\theta_{k}$, where $\mathbf{P}_{\mathbf{u}}[\theta_{k}]=\left[\mathbf{A}_{\theta_{k}}^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta_{k}}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right]^{-1}$. Writing out the IPLA update steps for this problem, we have
@@ 3.2 Solving the Inverse Problem for Linear Poisson PDE
$\theta_{k+1}$ $=\theta_{k}-\gamma\left(n_{u}+\frac{\theta_{k}-\mu_{\theta}}{\sigma_{\theta}^{2}}\right)-\frac{\gamma}{N}\sum_{n=1}^{N}\left(\mathbf{A}_{\theta_{k}}\mathbf{u}_{k}^{(n)}-\mathbf{b}\right)^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta_{k}}\mathbf{u}_{k}^{(n)}+\sqrt{\frac{2\gamma}{N}}\,\bm{\zeta}_{k+1}^{(0)}$ (29)
$\mathbf{u}_{k+1}^{(n)}$ $=(1-\gamma)\mathbf{u}_{k}^{(n)}+\gamma\mathbf{P}_{\mathbf{u}}[\theta_{k}]\left(\mathbf{A}_{\theta_{k}}^{\top}\mathbf{G}^{-1}\mathbf{b}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{y}\right)+\sqrt{2\gamma}\,\mathbf{P}_{\mathbf{u}}[\theta_{k}]^{1/2}\bm{\zeta}_{k+1}^{(n)}$ (30)
for all $n\in\left\{1,\dots,N\right\}$, with $\bm{\zeta}_{k}^{(0)}\sim\mathcal{N}(0,1)$, and $\{\bm{\zeta}_{k}^{(n)}\}_{n=1}^{N}$ as i.i.d. standard $n_{u}$-dimensional Gaussians. The factorisation $\mathbf{A}_{\theta}=e^{\theta}\mathbf{A}$ allows us to compute the gradient $\nabla_{\theta}\Psi_{\theta}^{y}$ for (29) analytically. For problems without such a factorisation, we could use automatic differentiation to compute $\frac{\partial}{\partial\theta}\mathbf{A}_{\theta}$ for use in the gradient update steps.
@@ 4.1 Convergence of elliptic PDE forcing estimate to MMAP estimator
Recall that we consider the setting in section 3.1 and in this case, we have the potential $\Psi^{y}_{\theta}(\mathbf{u},\mathbf{b})$ for the joint system $\mathbf{v}=(\mathbf{u},\mathbf{b})$ conditioned on the data $\mathbf{y}$. We can write the gradients of the potential w.r.t. the solution, $\mathbf{u}$, and parameters $\mathbf{b}$, are as follows
$\nabla_{\mathbf{u}}\Psi^{y}_{\theta}(\mathbf{u},\mathbf{b})$ $=\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\left(\mathbf{A}_{\theta}\mathbf{u}-\mathbf{b}\right)+\mathbf{H}^{\top}\mathbf{R}^{-1}\left(\mathbf{H}\mathbf{u}-\mathbf{y}\right),$ (31)
$\nabla_{\mathbf{b}}\Psi_{\theta}^{y}(\mathbf{u},\mathbf{b})$ $=-\mathbf{G}^{-1}\left(\mathbf{A}_{\theta}\mathbf{u}-\mathbf{b}\right)+\bm{\Sigma}^{-1}(\mathbf{b}-\bm{\mu}).$ (32)
For this problem we have a quadratic potential, and so the strong convexity and Lipschitz constants are determined by the minimum and maximum eigenvalues of the Hessian:
$\nabla^{2}\Psi_{\theta}^{y}=\begin{bmatrix}\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}&-\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\\ -\mathbf{G}^{-1}\mathbf{A}_{\theta}&\mathbf{G}^{-1}+\bm{\Sigma}^{-1}\end{bmatrix}.$ (33)
Provided the Hessian is positive-definite, we have $\mu>0$, and a unique minimiser. We can indeed prove this result for this particular problem.
@@
The potential $\Psi_{\theta}^{y}(\mathbf{u},\mathbf{b})$ for the linear-Poisson equation is $\mu$-strongly convex in $(\mathbf{u},\mathbf{b})$, with $\mu=\lambda_{\min}(\nabla^{2}\Psi_{\theta}^{y})>0$.
@@
We also show that without the prior, considering only the MMLE potential $\Phi_{\theta}^{y}$, we can only guarantee $\mu=\lambda_{\min}(\nabla^{2}\Phi_{\theta}^{y})\geq 0$ and therefore is only convex, not strongly convex. This motivates the use of the prior on the forcing function to ensure identifiability and guarantee a unique minimiser. For this linear case, we can determine this minimiser analytically and we derive this in Appendix C.
Theorem 1.
Assume $L:=\lambda_{\max}(\nabla^{2}\Phi_{\theta}^{y})<\infty$. Then, the parameter marginal $(\mathbf{b}_{k})_{k\geq 1}$ generated by the IPLA algorithm (24)–(25) with the step-size satisfying $\gamma\leq 2/(\mu+L)$ for estimating the forcing function in the linear Poisson equation converges to the MMAP estimator $\mathbf{b}^{\star}=\arg\max p(\mathbf{b}|\theta,\mathbf{y})$. In particular, we have the following convergence result
where $\nu_{0}$ is the initial distribution of the particle system so that $W_{2}(\nu_{0},\tilde{\pi})<\infty$ where $\tilde{\pi}$ is given in the proof, $\mu$ is the strong convexity constant given in Lemma 1.
Proof.
See Appendix B. ∎
We experimentally confirm the result in Theorem 1, by estimating the order of convergence of the IPLA estimate to the MMAP estimate in section 5.1. A few remarks to fully unpack this result are in order.
@@
Here, we note the linear dependence of the complexity on the latent dimension $n_{u}$, that is for lower latent dimension (i.e. for coarser FEM discretisation) the complexity will reduce, which may appear counter-intuitive. This is a result of bounding the error between the MMAP estimate $\mathbf{b}^{\star}$ and the IPLA estimate $\mathbf{b}_{k}$; with $n_{u}$ too small, the resulting MMAP estimate will not necessarily be close to the true forcing $\mathbf{b}$ due to the large FEM discretisation error incurred. We point readers towards the works [33, 39], which provide error bounds for statFEM between the true solution and the statFEM distributions, in terms of the FEM discretisation.
@@
Remark 1 makes it clear that the number of iterations $k$ to obtain $\varepsilon$-accuracy is strongly influenced by $\kappa$. It is thus crucial to control this number for ill-conditioned problems, as in our case. The algorithm in (27)–(28) is preconditioned and this can potentially improve our bound. Specifically, with a transformation of variables $\mathbf{w}\coloneqq\mathbf{P}^{-1/2}\mathbf{v}$, it can be shown via Itô’s formula that an equivalent SDE, with potential $\Psi^{y}_{\theta,\mathbf{P}}(\mathbf{w})\coloneqq\Psi_{\theta}^{y}(\mathbf{v})=\Psi^{y}(\mathbf{P}^{1/2}\mathbf{w})$ has the same invariant measure [4]. We can then analyse the Lipschitz and strong convexity constants in the transformed coordinate system, using the same method as in section 4.1 now that the SDE is of the same form. The Hessian for the transformed potential $\nabla_{\mathbf{w}}^{2}\Psi^{y}_{\theta,\mathbf{P}}(\mathbf{w})=\mathbf{P}^{1/2}\nabla_{\mathbf{v}}^{2}\Psi_{\theta}^{y}(\mathbf{v})\mathbf{P}^{1/2}$, and so the constants relating to the preconditioned system are
$\mu_{\mathbf{P}}=\lambda_{\mathrm{min}}\left(\mathbf{P}^{1/2}\nabla^{2}\Psi^{y}_{\theta}\mathbf{P}^{1/2}\right),\quad L_{\mathbf{P}}=\lambda_{\mathrm{max}}\left(\mathbf{P}^{1/2}\nabla^{2}\Psi_{\theta}^{y}\mathbf{P}^{1/2}\right).$ (35)
The bound in Theorem 1 then applies directly with the constants $\mu_{\mathbf{P}}$ and $L_{\mathbf{P}}$. Let $\kappa_{\mathbf{P}}=L_{\mathbf{P}}/\mu_{\mathbf{P}}$ be the condition number of the preconditioned system. Given Remark 1 and assuming that $\kappa_{\mathbf{P}}\ll\kappa$, preconditioning has two clear effects: First, the step-size $\gamma=\mathcal{O}(\varepsilon^{2}n_{u}^{-1}\kappa_{\mathbf{P}}^{-2})$ can now be bigger (within the restrictions provided in Theorem 1). Secondly, to obtain $\varepsilon$-accuracy, we may require significantly fewer steps, i.e., $k=\widetilde{\mathcal{O}}(n_{u}\kappa_{\mathbf{P}}^{2}\varepsilon^{-2})$. $\diamond$